# Dataset Decomposition: Faster LLM Training with Variable Sequence Length Curriculum

### v1

@@ Abstract
Large language models (LLMs) are commonly trained on datasets consisting of fixed-length token sequences. These datasets are created by randomly concatenating documents of various lengths and then chunking them into sequences of a predetermined target length. However, this method of concatenation can lead to cross-document attention within a sequence, which is neither a desirable learning signal nor computationally efficient. Additionally, training on long sequences becomes computationally prohibitive due to the quadratic cost of attention. In this study, we introduce dataset decomposition, a novel variable sequence length training technique, to tackle these challenges. We decompose a dataset into a union of buckets, each containing sequences of the same size extracted from a unique document. During training, we use variable sequence length and batch size, sampling simultaneously from all buckets with a curriculum. In contrast to the concat-and-chunk baseline, which incurs a fixed attention cost at every step of training, our proposed method incurs a penalty proportional to the actual document lengths at each step, resulting in significant savings in training time. We train an 8k context-length 1B model at the same cost as a 2k context-length model trained with the baseline approach. Experiments on a web-scale corpus demonstrate that our approach significantly enhances performance on standard language evaluations and long-context benchmarks, reaching target accuracy 3×3$\times$3 × faster compared to the baseline. Our method not only enables efficient pretraining on long sequences but also scales effectively with dataset size. Lastly, we shed light on a critical yet less studied aspect of training large language models: the distribution and curriculum of sequence lengths, which results in a non-negligible difference in performance.
@@ 1 Introduction
(a) Data scaling
(b) Model scaling
Figure 1: (a) OpenLM-1B models trained with baseline-8k and our method on up to 275B tokens. (b) OpenLM-1B/3B/7B models trained on 137B tokens. Accuracy and training speed gains are shown.
@@
The concat-and-chunk approach has several shortcomings. First, randomly concatenating documents can lead to the model attending to a context from an unrelated document to predict the next token. While well-trained models learn to avoid cross-document attention, this is not explicitly enforced, leading to potential spurious modeling. Second, the cross-document attention spends unnecessary computation on attending to unrelated tokens that do not facilitate learning. This is especially crucial due to the quadratic complexity of the attention mechanism. Third, even if a document is shorter than the target sequence length, it may still be broken into two chunks when they are at the boundary of two sequences. This results in significantly smaller average chunk lengths compared to the original document length average (see Fig. 3(a)), which hinders the model’s capability.
@@
Recent and concurrent works on LLM training try to improve the concat-and-chunk approach: document-masking [36] to resolve cross-document attention, best-fit packing [17] to reduce document chunking, and concatenating semantically related documents instead of randomly [51]. However, none of them address all three issues mentioned above together.
@@
In this work, we introduce dataset decomposition (DD), a novel approach to decompose data based on their length and train with variable sequence length (VSL) to address the above issues. DD decomposes a given dataset containing documents of variable lengths into a union of datasets/buckets, each with sequences of a fixed length. Specifically, a dataset is decomposed $\mathcal{D}$ into buckets $\cup_{i}\mathcal{D}_{i}$, where each bucket $\mathcal{D}_{i}$ contains sequences of length $2^{i}$, each extracted from a unique document. During training with VSL, at every step of the optimization process, we sample $i$ (based on a curriculum) to form a batch with $b/2^{i}$ sequences from the bucket $\mathcal{D}_{i}$, which keeps the total number of tokens in a batch constant at $=b$, regardless of which $\mathcal{D}_{i}$ is sampled.
@@
This approach gives us several advantages and resolves the aforementioned issues of the concat-and-chunk method. First, DD is simple and has negligible computational overhead during the data preparation stage, making it easy to scale to large datasets. Second, tokens in each sequence from every bucket are ensured to be from the same document by construction, which avoids cross-document attention. Furthermore, we have access to the sequence length distribution (an auxiliary prior knowledge) which can be used to create different mixtures/curricula for training. Finally, our VSL training strategy accelerates training time: the latency for one optimization step is less when sampling from $\mathcal{D}_{i}$ with smaller $i$ (due to attention’s quadratic complexity). Following is a summary of our contributions:
@@
  • We perform large-scale experimentation using different models, datasets, and evaluation tasks to demonstrate the efficacy of the proposed method. We show (see Fig. 1) significant gains in data efficiency (>2×>2$\times$> 2 ×) and compute efficiency (11% to 45%), resulting in combined LLM pretraining acceleration of up to 3×3$\times$3 × (time to reach certain accuracy compared to baseline).
@@ 3 Experiments and analysis
  • Multi-Document Question Answering (MDQA): We follow the exact setup as in [32], where for each question from NaturalQuestions-Open [27, 26], $r$ Wikipedia documents are retrieved such that one of them has the answer to the question, and the other $r-1$ documents are distractors. We report MDQA-10, MDQA-20, and MDQA-30 accuracy corresponding to $r=10,20$, and 30, respectively. For each query, we evaluate the model by changing the location of the target document among distractors and report the averaged accuracy.
  • TOEFL: This dataset is a multiple-choice question answering dataset from [5]. The dataset contains QA pairs for 15 longest lectures in [57, 11]. Only one of the choices is the correct response. We estimate the correct choice by picking the choice with the lowest mean log probability value.
  • QuALITY: This dataset is a multiple-choice question answering dataset from [5]. The dataset contains a long passage for context, followed by a question with multiple choices. Only one of the choices is the correct response. We estimate the correct choice by picking the choice with the lowest mean log probability value.
Figure 4: Average time for one optimization step ($b=8\times 8192$ tokens) on an 8$\times$×H100 node with FSDP and FlashAttention2 for different context lengths.
@@ 3.1 Training efficiency
We first verify that VSL training enables a higher throughput than the baseline concat-and-chunk method. We enumerate model sizes (OpenLM-1B/3B/7B) and different context lengths ($2^{6}$ to $2^{13}$) and measure the time to train 100 batches with a fixed global batchsize of $b=8\times 8192$ distributed over $8$ GPUs in a single node. We repeat this 5 times and report the average time per optimization step in Fig. 4 (with STD mostly $<1$ms). See Section C.1 for additional results with different batchsizes $b$. For each model, we highlight the training time overhead (due to attention’s quadratic complexity with an optimized FlashAttention2 kernel [15]) when training with 8192 context lengths compared to 64 context lengths: +35%, +88%, and +23% for OpenLM-1B, -3B [^2] OpenLM-3B has 32 heads (×$\times 2$ that of OpenLM-1B), and a per-head dimension of 2560/32=80, not suitable for FlashAttention. This makes attention a significant bottleneck for this model., and -7B, respectively. Training overhead grows for longer context lengths (see Fig. 7 for results up to 16k context length).
@@ 3.3 Data mixture
A key benefit of dataset decomposition is access to and control over sequence length distribution. We form datasets with different mixtures of sequence lengths and explore the performance of a model trained on each mixture. Table 1 shows the results. For all experiments, the total seen tokens and hyperparameters are fixed, and only the distribution over sequence length is changed. First, we observe that mixtures with small average context length (we provide the exact definition in Appendix F) perform poorly on MDQA, which requires long context understanding. Larger average context length also correlates positively with performance on reading comprehension tasks, consistent with our observation in Fig. 5(a), but comes at the cost of a longer training step.
Furthermore, “1k-only”, that is training using only the best sequence length ($=1024$) from the study in Section 3.2 results in good performance on regular evaluations, especially for language understanding and world knowledge tasks, but is poor for long context tasks. Finally, we observe that “natural” mixture, that is aligned with the distribution resulting from dataset decomposition (see Fig. 3(b)), obtains near-optimal performance on both regular and MDQA tasks, demonstrating the scalability of the proposed approach to large datasets.
@@ 3.4 Length-based curriculum
We can think of short sequences as being "easier" compared to longer ones; hence motivating a curriculum learning [7, 18] that prioritizes short sequences. A similar idea (training with image resolutions from low to high) is explored in vision to train CLIP [45] models more efficiently [31]. In VSL, we can easily implement curriculum learning through sampling designs. At every optimization step, we sample without replacement a batch with $b$ tokens from bucket $\mathcal{D}_{i}$ with probability $p_{i}$. If a bucket is empty, we exclude it from sampling. We study different curricula for the "$\geq 256$" mixture (with an equal number of tokens in $\mathcal{D}_{8},\ldots,\mathcal{D}_{13}$). Results are shown in Table 2. For each curriculum, we determine the odds of picking a batch from each bucket ($=p_{i}$’s when normalized). We consider curricula that shift from short to long sequences at different paces controlled by $p_{i}$’s changing linearly, with powers of 2, and with powers of 100 between buckets.
Due to the presence of other hyperparameter schedules during the course of training (e.g., learning rate and weight decay), a curriculum on length may result in a potential implicit bias. For example, if we only see long sequences toward the end of training, long sequence learning occurs only when the learning rate is too small. To address this potential issue, we also explore cyclic curricula, where a curriculum is applied in cycles (similar to cyclic learning rate schedules [52]).
Our results show that the cyclic "Grow-P2" curriculum is near optimal with different metrics. An additional benefit of curriculum is training stability. [30] noticed that long sequences contribute to extreme gradient variance, especially at the beginning of training, resulting in instability. We also observe (see Appendix E) that our proposed approach with curriculum results in more stable training dynamics, thus enabling more efficient training with larger batch sizes and learning rates.
@@ 3.5 Scaling
We demonstrate the efficacy of our proposed method on another large-scale dataset. We use an improved variant of the RefinedWeb [42] data pipeline to filter CommonCrawl [2] and construct a pretraining dataset. We train models with different numbers of parameters: OpenLM-160M, -410M, and -1B, for a total of 137B tokens. We compare the baseline with a $DD_{\geq 256}$ mixture trained with the "Grow-P2" curriculum with 8 cycles. Results are reported in Table 3, demonstrating significant accuracy and training efficiency gains.
@@ 3.6 Comparison with state-of-the-art
Document masking improves the baseline on regular evaluations from $51.5$ to $52.4$ by preventing cross-document attention. However, [60] demonstrate that including concatenated unrelated documents can still enhance long-context metrics compared to training solely with shorter sequences. Therefore, DM experiences a slight decline in long-context evaluations, dropping from $27.5$ to $27.1$. Best-fit packing [17] with DM addresses this by arranging documents more effectively, thus reducing chunking/truncation and resulting in longer document segments in each training sequence. This approach enhances performance in both regular and long-context evaluations.
On the other hand, ICLM [51] proposes document sorting based on content similarity. Although the benefits of ICLM with large-scale Common Crawl data (used in our experiments) are marginal in regular evaluation, it significantly boosts long-context evaluation metrics from $27.5$ to $28.7$ by constructing long data sequences through the concatenation of similar documents. However, the similarity finding step proposed by ICLM is resource-intensive at scale [^4] Processing 400B tokens with the official repository required over a week using 96 CPUs and 16 GPUs.. Lastly, our proposed method avoids cross-document attention to unrelated content, maintains coherent long sequences, and benefits from a length-based curriculum, effectively improving performance in both regular and long-context evaluations.
@@ 4 Related works
Related to our study on sequence length bias, [58] shows the importance of train-vs-test time distribution shift from a sequence length perspective on a string editing task. [6, 62, 24, 33] highlight the challenge of generalizing to lengths beyond what the model has seen during training and discuss the importance of positional encoding. Several works [38, 63, 23, 60, 10, 43, 44, 49] address enabling LLM inference with long context (see [41] for an overview). These approaches are orthogonal to our contribution and can be applied post-pretraining to adapt to longer lengths.

### v2

@@ Abstract
Large language models (LLMs) are commonly trained on datasets consisting of fixed-length token sequences. These datasets are created by randomly concatenating documents of various lengths and then chunking them into sequences of a predetermined target length (concat-and-chunk). Recent attention implementations mask cross-document attention, reducing the effective length of a chunk of tokens. Additionally, training on long sequences becomes computationally prohibitive due to the quadratic cost of attention. In this study, we introduce dataset decomposition, a novel variable sequence length training technique, to tackle these challenges. We decompose a dataset into a union of buckets, each containing sequences of the same size extracted from a unique document. During training, we use variable sequence length and batch-size, sampling simultaneously from all buckets with a curriculum. In contrast to the concat-and-chunk baseline, which incurs a fixed attention cost at every step of training, our proposed method incurs a computational cost proportional to the actual document lengths at each step, resulting in significant savings in training time. We train an 8k context-length 1B model at the same cost as a 2k context-length model trained with the baseline approach. Experiments on a web-scale corpus demonstrate that our approach significantly enhances performance on standard language evaluations and long-context benchmarks, reaching target accuracy with up to 6×6$\times$6 × faster training compared to the baseline. Our method not only enables efficient pretraining on long sequences but also scales effectively with dataset size. Lastly, we shed light on a critical yet less studied aspect of training large language models: the distribution and curriculum of sequence lengths, which results in a non-negligible difference in performance. [^1] Code to be available at https://github.com/apple/ml-dataset-decomposition.
[…]
@@ 1 Introduction
(a) Data Efficiency
(b) Computational Efficiency
Figure 1: Scaling the training of the OpenLM-410M model to 1.1 trillion tokens on the RefinedWeb dataset. Note that each point on the figure represents a separate training from scratch (not different checkpoints of a single run). For the largest run, some tokens are seen more than once (the dataset has approximately 525 billion tokens). For dataset decomposition, we use the Grow-P2 curriculum with 8 cycles. Maximum context length is 8192 and all hyper-parameters are the same for DD and the baseline. (a) Regular metrics average when training with the baseline method and the proposed method. We observe more than 4×4$\times$4 × data efficiency compared to the baseline. Also, even at 1.1 trillion tokens, DD has a +2.4 accuracy improvement compared to the baseline (which has a plateauing accuracy curve even on a logarithmic x-axis). (b) Comparison of model average accuracy versus training cost (GPU-hours). DD reaches the best accuracy of the baseline more than 6×6$\times$6 × faster. This is the combined effect of DD accuracy and speed gains.
@@
The concat-and-chunk approach has several shortcomings. First, randomly concatenating documents can lead to the model attending to a context from an unrelated document to predict the next token. While well-trained models learn to avoid cross-document attention, this is not explicitly enforced, leading to potential spurious modeling. Second, the cross-document attention spends unnecessary computation on attending to unrelated tokens that do not facilitate learning. This is especially crucial due to the quadratic complexity of the attention mechanism. Even with an implementation of attention that supports cross-document attention masking, the computational cost for each optimization step would be bottlenecked by the longest document in the global batch, leading to significant under-utilization of devices with shorter documents. Third, even if a document is shorter than the target sequence length, it may still be broken into two chunks when they are at the boundary of two sequences. This results in significantly smaller average chunk lengths compared to the original document length average (see Fig. 3(a)), which hinders the model’s capability.
@@
Recent and concurrent works on LLM training try to improve the concat-and-chunk approach: document-masking is possible with recent implementation of attention [29] as adopted in some recent pre-training recipes [39], best-fit packing [17] to reduce document chunking, and concatenating semantically related documents instead of randomly [55]. However, none of them address all three issues mentioned above together.
@@
In this work, we introduce dataset decomposition (DD), a novel approach to decompose data based on their length and train with variable sequence length (VSL) and length-based curriculum to address the above issues. We obtain significant both significant accuracy improvement and straining speed-up as shown in Fig. 1. DD decomposes a given dataset containing documents of variable lengths into a union of datasets/buckets, each with sequences of a fixed length. Specifically, a dataset $\mathcal{D}$ is decomposed into buckets $\cup_{i}\mathcal{D}_{i}$, where each bucket $\mathcal{D}_{i}$ contains sequences of length $2^{i}$, each extracted from a unique document. During training with VSL, at every step of the optimization process, we sample $i$ (based on a curriculum) to form a batch with $b/2^{i}$ sequences from the bucket $\mathcal{D}_{i}$, which keeps the total number of tokens in a batch constant ($2^{i}\times b/2^{i}=b$), regardless of which $\mathcal{D}_{i}$ is sampled.
@@
This approach gives us several advantages and resolves the aforementioned issues of the concat-and-chunk method. First, DD is simple and has negligible computational overhead during the data preparation stage, making it easy to scale to large datasets. Second, tokens in each sequence are ensured to be from the same document by construction, which avoids cross-document attention. Furthermore, we have access to the sequence length distribution (an auxiliary prior knowledge) which can be used to create different mixtures/curricula for training. Finally, our VSL training strategy accelerates training time: the latency for one optimization step is less when sampling from $\mathcal{D}_{i}$ with smaller $i$ (due to attention’s quadratic complexity). Following is a summary of our contributions:
@@
  • We perform large-scale experimentation using different models, datasets, and evaluation tasks to demonstrate the efficacy of the proposed method. We show (see Fig. 1) significant gains in data efficiency (>4×>4$\times$> 4 ×) and compute efficiency (11% to 45%), resulting in combined LLM pretraining acceleration of up to 6×6$\times$6 × (time to reach certain accuracy compared to baseline).
@@ 3 Experiments and analysis
Figure 4: (a) Average time for one optimization step ($b=8\times 8192$ tokens) on an 8$\times$×H100 node with FSDP and FlashAttention2 for different context lengths. (b) OpenLM-1B/3B/7B models trained on 137B tokens. Accuracy and training speed gains are shown.
@@
  • Multi-Document Question Answering (MDQA): We follow the exact setup as in Liu et al. [35], where for each question from NaturalQuestions-Open [28, 27], $r$ Wikipedia documents are retrieved such that one of them has the answer to the question, and the other $r-1$ documents are distractors. We report MDQA-10, MDQA-20, and MDQA-30 accuracy corresponding to $r=10,20$, and 30, respectively. For each query, we evaluate the model by changing the location of the target document among distractors and report the averaged accuracy.
  • TOEFL: This dataset is a multiple-choice question answering dataset from An et al. [5]. The dataset contains QA pairs for 15 longest lectures in Tseng et al. [61], Chung et al. [11]. Only one of the choices is the correct response. We estimate the correct choice by picking the choice with the lowest mean log probability value.
  • QuALITY: This dataset is a multiple-choice question answering dataset from An et al. [5]. The dataset contains a long passage for context, followed by a question with multiple choices. Only one of the choices is the correct response. We estimate the correct choice by picking the choice with the lowest mean log probability value.
@@ 3.1 Training efficiency
We first verify that VSL training enables a higher throughput than the baseline concat-and-chunk method. We enumerate model sizes (OpenLM-1B/3B/7B) and different context lengths ($2^{6}$ to $2^{13}$) and measure the time to train 100 batches with a fixed global batchsize of $b=8\times 8192$ distributed over $8$ GPUs in a single node. We repeat this 5 times and report the average time per optimization step in Fig. 4(a) (with STD mostly $<1$ms). See Section C.1 for additional results with different batchsizes $b$. For each model, we highlight the training time overhead (due to attention’s quadratic complexity with an optimized FlashAttention2 kernel [15]) when training with 8192 context lengths compared to 64 context lengths: +35%, +88%, and +23% for OpenLM-1B, -3B [^3] OpenLM-3B has 32 heads (×$\times 2$ that of OpenLM-1B), and a per-head dimension of 2560/32=80, not suitable for FlashAttention. This makes attention a significant bottleneck for this model., and -7B, respectively. Training overhead grows for longer context lengths (see Fig. 7 for results up to 16k context length).
@@ 3.3 Data mixture
A key benefit of dataset decomposition is access to and control over sequence length distribution. We form datasets with different mixtures of sequence lengths and explore the performance of a model trained on each mixture. Table 1 shows the results. For all experiments, the total seen tokens and hyperparameters are fixed, and only the distribution over sequence length is changed. First, we observe that mixtures with small average context length (we provide the exact definition in Appendix F) perform poorly on MDQA, which requires long context understanding. For example, as for “1k-only”, $\leq$2k”, and “Mid” distributions that do not include long sequences from $\mathcal{D}_{12}$ and $\mathcal{D}_{13}$. Larger average context length (e.g., as in $\geq$1k”) also correlates positively with performance on reading comprehension tasks, consistent with our observation in Fig. 5(a), but comes at the cost of a longer training step time.
Furthermore, “1k-only”, that is training using only the best sequence length ($=1024$) from the study in Section 3.2 results in good performance on regular evaluations, especially for language understanding and world knowledge tasks, but is poor for long context tasks. Finally, we observe that “natural” mixture, that is aligned with the distribution resulting from dataset decomposition (see Fig. 3(b)), obtains near-optimal performance on both regular and MDQA tasks, demonstrating the scalability of the proposed approach to large datasets without a need for intervention on the natural underlying length distribution.
@@ 3.4 Length-based curriculum
We can think of short sequences as being "easier" compared to longer ones; hence motivating a curriculum learning [7, 18] that prioritizes short sequences. A similar idea (training with image resolutions from low to high) is explored in vision to train CLIP [49] models more efficiently [33]. In VSL, we can easily implement curriculum learning through sampling designs. At every optimization step, we sample without replacement a batch with $b$ tokens from bucket $\mathcal{D}_{i}$ with probability $p_{i}$. If a bucket is empty, we exclude it from sampling. We study different curricula for the "$\geq 256$" mixture (with an equal number of tokens in $\mathcal{D}_{8},\ldots,\mathcal{D}_{13}$). Results are shown in Table 2. For each curriculum, we determine the odds of picking a batch from each bucket ($=p_{i}$’s when normalized). Details of our length-based sampling and curriculum are provided in Algorithm 1. We consider curricula that shift from short to long sequences at different paces controlled by $p_{i}$’s changing linearly, with powers of 2, and with powers of 100 between buckets.
Due to the presence of other hyperparameter schedules during the course of training (e.g., learning rate and weight decay), a curriculum on length may result in a potential implicit bias. For example, if we only see long sequences toward the end of training, long sequence learning occurs only when the learning rate is too small. To address this potential issue, we also explore cyclic curricula, where a curriculum is applied in cycles similar to cyclic learning rate schedules [56] as shown in Fig. 6. Note that when we train on a sequence of length $l$, we have $l$ next-token prediction losses (applied in parallel) with context lengths $10,1,\ldots,l-$ 1. This already implies some mixing: when training on a “hard” example (i.e., a long sequence), we also include “easy” examples (its shorter sub-sequences). Therefore, even towards the end of each cycle, we still have some losses with short contexts.
Our results show that the cyclic "Grow-P2" curriculum is near optimal with different metrics. An additional benefit of curriculum is training stability. Li et al. [31] noticed that long sequences contribute to extreme gradient variance, especially at the beginning of training, resulting in instability. We also observe (see Appendix E) that our proposed approach with curriculum results in more stable training dynamics, thus enabling more efficient training with larger batch sizes and learning rates.
@@ 3.5 Scaling
We demonstrate the efficacy of our proposed method on another large-scale dataset, DataComp-LM [32]. We train models with different numbers of parameters: OpenLM-160M, -410M, and -1B, for a total of 137B tokens. We compare the baseline with a $DD_{\geq 256}$ mixture trained with the "Grow-P2" curriculum with 8 cycles. Results are reported in Table 3, demonstrating significant accuracy and training efficiency gains.
@@ 3.6 Comparison with state-of-the-art
Pre-training context length is an important factor in determining a model’s long-context performance. We empirically validate this in the results shown in Fig. 5(a), where models trained on longer sequences perform better on multi-document QA. Our proposed method has an average context length (as defined in Eq. 2) of 1,344 for the RefinedWeb dataset, compared to 930 for the baseline (see Fig. 3(c)) and 1,064 when packing [17] is applied. This explains why the dataset decomposition mixture, even without any length-based curriculum (the first row in Table 2), outperforms Baseline-8k-DM and Pack-8k+DM (second and third rows in Table 5). Here, DM refers to applying document masking during training to avoid cross-document attention.
Document masking improves the baseline on regular evaluations from $51.5$ to $52.4$ by preventing cross-document attention. However, Xiong et al. [64] demonstrate that including concatenated unrelated documents can still enhance long-context metrics compared to training solely with shorter sequences. Therefore, DM experiences a slight decline in long-context evaluations, dropping from $27.5$ to $27.1$. Baseline-8k multi-document QA performance is even slightly better than our proposed dataset decomposition mixture when used without length-based curriculum (the first row in Table 2).
In-context pre-training LMs (ICLM) [55] proposes document sorting based on content similarity. Although the benefits of ICLM with large-scale Common Crawl data (used in our experiments) are marginal in regular evaluation, we observe that ICLM results in slightly better multi-document QA performance than Baseline-8k when 30 documents are in the context compared with Baseline-8k (22.0% vs. 20.5%). The average long-context metric boosts from $27.5$ for Baseline-8k to $28.7$ for ICLM. However, the similarity finding step proposed by ICLM is resource-intensive at scale [^5] Processing 400B tokens with the official repository required over a week using 96 CPUs and 16 GPUs..
Finally, as shown in in Table 2 our proposed cyclic length-based curriculum, for example, Grow-P2 with 8 cycles, results in a significant improvement in the model’s long-context capability. Our proposed method avoids cross-document attention to unrelated content, maintains coherent long sequences, and benefits from a length-based curriculum, effectively improving performance in both regular and long-context evaluations compared to all baselines. We further summarize long-context performance of different methods discussed above in Table 6.
@@ 4 Related works
Method Doc Masking Average Context Docs in a Sequence Curr. MDQA-30 (%)
Baseline 930 Mult-random 16.0
Pack-8k+DM [17] 1064 Mult-packing 16.9
DD-Uniform 1344 Single 19.6
Baseline 4096 Mult-random 20.5
ICLM [55] 4096 Mult-semantic 22.0
DD-Grow-P2 N/A 1344 Single 24.6
Table 6: Summary of long-context performance for different methods from Table 2 and Table 5.
@@
Related to our study on sequence length bias, [62] shows the importance of train-vs-test time distribution shift from a sequence length perspective on a string editing task. [6, 66, 25, 36] highlight the challenge of generalizing to lengths beyond what the model has seen during training and discuss the importance of positional encoding. Several works [41, 67, 23, 64, 10, 47, 48, 53, 34] address enabling LLM inference with long context (see [45] for an overview). These approaches are orthogonal to our contribution and can be applied post-pretraining to adapt to longer lengths. GrowLength [24] proposes accelerating LLM pretraining by progressively growing context length using the baseline sequence formation method, but does not show results on LLMs. Similarly, increasing sequence length has been shown in BERT model training [42] to improve compute efficiency.